OpenStudy (anonymous):

An experiment consists of rolling a fair die and flipping a fair coin. What is the probability of getting an even number and a head? a3/4 b0 c1/6 d1/4

OpenStudy (zarkon):

p(even number and a head)=P(even number)*P(a head)

OpenStudy (anonymous):

what?

OpenStudy (campbell_st):

P(even No) = 1/2 P (head) = 1/2 P(even and Head) = 1/2 x 1/2 = 1/4